# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v1

@@ 4.1 The localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$
(see [13, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.71.
(We also note that Ryan–Smith’s localization for $\mathsf{PP}$ uses both conjuncts: $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$; see [13, Prop. 3.17] and Question 3.18 therein, and see [13, Props. 3.22 and 3.23] for counterexamples showing the conjuncts cannot in general be omitted.)
@@
Fix $p\in\mathbb{Q}_{f}$ and $x\in X$ with $x\notin\operatorname{dom}(p)$. Choose any $y\in Y$ with $f(y)=x$ (exists since $f$ is surjective). We claim $y\notin\operatorname{ran}(p)$: if $y=p(x^{\prime})$ for some $x^{\prime}\in\operatorname{dom}(p)$ then $x=f(y)=f(p(x^{\prime}))=x^{\prime}$, contradicting $x\notin\operatorname{dom}(p)$. Thus $q:=p\cup\{(x,y)\}$ is still a partial function, still injective, and still satisfies $f(q(x))=x$; hence $q\in\mathbb{Q}_{f}$ and $q\leq p$ with $x\in\operatorname{dom}(q)$. ∎
@@ 4.2 The $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{f}$
In $\mathsf{ZF}$ the ordinal $\aleph^{*}(S)$ exists for every set $S$ (it is the Lindenbaum number of $S$); see e.g. Jech [4, §(Hartogs/Lindenbaum)].
@@
Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [13, Cor. 3.3]. For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [13, Prop. 3.2]. Thus, in the presence of $\mathsf{SVC}(S)$ it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)$.
@@
$\mathbb{R}_{f}$ is Countably closed.
@@
Finally, $\operatorname{dom}(r)=\bigcup_{n<\omega}\operatorname{dom}(r_{n})\subseteq\lambda$ is a countable union of countable sets of ordinals, hence countable in $M$ (this uses only $\mathsf{ZF}$ since $\lambda$ is well-ordered and each countable subset of an ordinal admits a canonical increasing enumeration). Therefore $r\in\mathbb{R}_{f}$, and clearly $r\leq r_{n}$ for all $n$. ∎
@@ 5.1 Package symmetric systems at a stage
At limit stages, $\omega_{1}$-completeness (and normality) of the limit filter is provided by the iteration machinery (API item (d), cf. [3, Thm. 5.5]).
@@ 5.2 Iteration-level diagonal automorphisms
This is exactly the successor-step group structure supplied by the countable-support symmetric iteration framework (Remark 2.1, cf. [3, Def. 3.1]). Concretely:
@@
(i) This is part of the inverse/direct system structure of the iteration groups (cf. [3, Def. 3.1]) and the fact that $\pi^{0}_{\lambda}$ is defined stagewise and respects composition.
@@
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are (by definition) coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf. [3, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection.
@@
Limit $\eta$: Define $g_{\eta}$ to be the direct-limit element of $\mathscr{G}_{\eta}$ determined by the family $\langle g_{\beta}:\beta<\eta\rangle$, which is coherent by the inductive invariant above. By the limit-stage group definition ([3, Def. 3.1]), $g_{\eta}\in\mathscr{G}_{\eta}$.
By the uniqueness property of the inverse/direct limit presentation: elements of $\mathscr{G}_{\eta}$ are determined by their projections to all earlier stages ([3, Def. 3.1]). Since $g_{\eta}$ and $\widehat{\pi}^{\,D_{\eta}}$ have identical restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$ (namely, $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$), we conclude $g_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
@@ 5.3 The modified limit filter
By Lemma 5.52, the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ is a symmetric system in the sense of [4, Ch. 15]. By the symmetric extension theorem [4, Lemma. 15.51], $M_{\lambda}\models\mathsf{ZF}$. ∎
@@ 5.4 Bookkeeping: names versus interpretations
    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.19 and 4.29), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
@@ 5.5 The symmetric iteration (black box)
We now invoke the iteration framework via Remark 2.1. For the global conclusion (a final model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$), we take a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$. Concretely, working in the metatheory, we form an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [3, §2.1], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
@@
At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group/filter as provided by the symmetric-iteration framework (API items (b)–(e), especially filter generation/completeness at limits, [3, Thm. 5.5]). We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@ 5.6 What the iteration yields
Choose a hereditarily symmetric name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$. Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{f}$ (API item (e)), and let $\alpha:=\sup(I)+1$. Then $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name and $f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}$. Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name, its interpretation is unchanged by further forcing: $\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}$, so $f=f_{\alpha}$ in $\mathcal{M}$.
@@
Remark 5.63 (Countable‐support names at limits of cofinality $\omega$).
For limits $\lambda$ of countable cofinality, a name may have countable support cofinal in $\lambda$. The iteration framework of [3] handles this by block‐factorization and limit‐filter completeness (see [3, Rem. 3.9, Thm. 5.5]); our argument implicitly relies on that machinery.
@@
By equivariance of name interpretation ([4, Lemma 14.37]):
@@
By Lemma 5.6, each orbit package $\mathbb{S}_{[f]}$ used at stage $\alpha$ is $\omega_{1}$-closed (and hence adds no new $\omega$-sequences of ordinals), and the finite-support product $\dot{\mathbb{S}}_{\alpha}$ is again $\omega_{1}$-closed for the same reason. By Lemma 5.16, the relevant stage filter is $\omega_{1}$-complete. Therefore the $\mathsf{DC}$-preservation theorem for symmetric extensions applies (API item (c), cf. [3, Thm. 3.20]), yielding $M_{\alpha+1}\models\mathsf{DC}$.
At limit stages, $\mathsf{DC}$ is preserved by the countable-support symmetric iteration machinery (API item (c), and the limit-filter generation/completeness from API item (d); cf. [3, Thm. 3.20] and [3, Thm. 5.5] for the limit construction).
@@
By Proposition 5.62, $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$, hence $\mathcal{M}\models\mathsf{PP}\!\restriction T$. By Proposition 5.71, $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. By Lemma 5.70, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ for $T=\mathcal{P}(S)$, so the hypotheses of the Ryan–Smith localization theorem (Theorem 4.5, i.e., [13, Prop. 3.17] applied with Ryan–Smith’s parameter $S$ taken to be our $T$) hold in $\mathcal{M}$. Therefore $\mathcal{M}\models\mathsf{PP}$.

### v3

@@ 4.1 The localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$
(see [17, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.71.
(We also note that Ryan–Smith’s localization for $\mathsf{PP}$ uses both conjuncts: $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$; see [17, Prop. 3.17] and Question 3.18 therein, and see [17, Props. 3.22 and 3.23] for counterexamples showing the conjuncts cannot in general be omitted.)
@@
Fix $p\in\mathbb{Q}_{f}$ and $x\in X$ with $x\notin\operatorname{dom}(p)$. Let $y\in Y$ be such that $f(y)=x$ (exists since $f$ is surjective). We claim $y\notin\operatorname{ran}(p)$: if $y=p(x^{\prime})$ for some $x^{\prime}\in\operatorname{dom}(p)$ then $x=f(y)=f(p(x^{\prime}))=x^{\prime}$, contradicting $x\notin\operatorname{dom}(p)$. Thus $q:=p\cup\{(x,y)\}$ is still a partial function, still injective, and still satisfies $f(q(x))=x$; hence $q\in\mathbb{Q}_{f}$ and $q\leq p$ with $x\in\operatorname{dom}(q)$. ∎
@@ 4.2 The $\mathsf{AC}_{\mathsf{WO}}$-package forcing $\mathbb{R}_{f}$
In $\mathsf{ZF}$ the ordinal $\aleph^{*}(S)$ exists for every set $S$ (it is the Lindenbaum number of $S$); see e.g. Jech [8, §(Hartogs/Lindenbaum)].
@@
Assuming $\mathsf{SVC}(S)$, Ryan–Smith shows that $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the bounded scheme $\mathsf{AC}_{<\aleph^{*}(S)}$ (more precisely, $\mathsf{AC}_{<\aleph^{*}(S)}(S)$ in his notation); see [17, Cor. 3.3]. For fixed $\lambda$, the corresponding “splitting” reformulations of $\mathsf{AC}_{\lambda}$ appear in [17, Prop. 3.2]. Thus, in the presence of $\mathsf{SVC}(S)$ it suffices to schedule splittings only for surjections onto ordinals $\lambda<\aleph^{*}(S)$.
@@
Assume $M\models\mathsf{ZF}+\mathsf{DC}$. Then $\mathbb{R}_{f}$ is countably closed.
@@
Finally, $\operatorname{dom}(r)=\bigcup_{n<\omega}\operatorname{dom}(r_{n})$ is a countable union of countable sets. Since $M\models\mathsf{DC}$ (hence $M\models AC_{\omega}$), $\operatorname{dom}(r)$ is countable in $M$. Therefore $r\in\mathbb{R}_{f}$, and clearly $r\leq r_{n}$ for all $n$. ∎
Remark 4.28 (Closure depends on $\mathsf{DC}$).
The $\omega_{1}$-closure of $\mathbb{R}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $AC_{\omega}$), via “countable union of countable sets is countable”. In bare $\mathsf{ZF}$ this need not hold.
@@ 5.1 Package symmetric systems at a stage
At limit stages, $\omega_{1}$-completeness (and normality) of the limit filter is provided by the iteration machinery (API item (d), cf. [5, Thm. 5.5]).
@@ 5.2 Iteration-level diagonal automorphisms
This is exactly the successor-step group structure supplied by the countable-support symmetric iteration framework (Remark 2.1, cf. [5, Def. 3.1]). Concretely:
@@
(i) This is part of the inverse/direct system structure of the iteration groups (cf. [5, Def. 3.1]) and the fact that $\pi^{0}_{\lambda}$ is defined stagewise and respects composition.
@@
At a limit stage $\eta$, elements of $\mathscr{G}_{\eta}$ are (by definition) coherent direct-limit elements whose restrictions to earlier stages determine them uniquely (cf. [5, Def. 3.1]). The induced bijection on $\mathrm{Coords}_{\eta}=\bigcup_{\beta<\eta}\mathrm{Coords}_{\beta}$ is obtained by taking the union of the stagewise bijections, and is again a bijection.
@@
Limit $\eta$: Define $g_{\eta}$ to be the direct-limit element of $\mathscr{G}_{\eta}$ determined by the family $\langle g_{\beta}:\beta<\eta\rangle$, which is coherent by the inductive invariant above. By the limit-stage group definition ([5, Def. 3.1]), $g_{\eta}\in\mathscr{G}_{\eta}$.
By the uniqueness property of the inverse/direct limit presentation: elements of $\mathscr{G}_{\eta}$ are determined by their projections to all earlier stages ([5, Def. 3.1]). Since $g_{\eta}$ and $\widehat{\pi}^{\,D_{\eta}}$ have identical restrictions to each $\mathbb{P}_{\beta}$ for $\beta<\eta$ (namely, $g_{\beta}=\widehat{\pi}^{\,D_{\beta}}$), we conclude $g_{\eta}=\widehat{\pi}^{\,D_{\eta}}$.
@@ 5.3 The modified limit filter
By Lemma 5.52, the triple $(\mathbb{P}_{\lambda},\mathscr{G}_{\lambda},\tilde{\mathscr{F}}^{*}_{\lambda})$ is a symmetric system in the sense of [8, Ch. 15]. By the symmetric extension theorem [8, Lemma. 15.51], $M_{\lambda}\models\mathsf{ZF}$. ∎
@@ 5.4 Bookkeeping: names versus interpretations
    The seed parameter $S=A^{\omega}$ is preserved at successor stages: our package forcings add no new reals (Corollaries 4.19 and 4.30), hence add no new $\omega$-sequences of Cohen reals and therefore no new elements of $S$.
@@ 5.5 The symmetric iteration (black box)
We now invoke the iteration framework via Remark 2.1. For the global conclusion (a final model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$), we take a class-length countable-support symmetric iteration of length $\mathrm{Ord}$ over $\mathcal{N}$. Concretely, working in the metatheory, we form an $\mathrm{Ord}$-indexed symmetric iteration $\langle\mathbb{P}_{\alpha},\mathscr{G}_{\alpha},\tilde{\mathscr{F}}^{*}_{\alpha}:\alpha\in\mathrm{Ord}\rangle$ as in [5, §2.1], with $\mathbb{P}_{0}=\operatorname{Add}(\omega,\omega_{1})$ and intermediate symmetric models $M_{\alpha}$.
@@
At limit stages $\lambda$, we take the countable-support limit forcing $\mathbb{P}_{\lambda}$ and the corresponding limit group/filter as provided by the symmetric-iteration framework (API items (b)–(e), especially filter generation/completeness at limits, [5, Thm. 5.5]). We write $\mathcal{M}$ for the final symmetric model of the $\mathrm{Ord}$-length iteration.
@@ 5.6 What the iteration yields
Choose a hereditarily symmetric name $\dot{f}$ such that $f=\dot{f}^{\,G_{\infty}}$. Let $I\subseteq\mathrm{Ord}$ be the countable stage support of $\dot{f}$ (API item (e)), and let $\alpha:=\sup(I)+1$. Then $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name and $f_{\alpha}:=\dot{f}^{\,G_{\alpha}}\in M_{\alpha}$. Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$-name, its interpretation is unchanged by further forcing: $\dot{f}^{\,G_{\infty}}=\dot{f}^{\,G_{\alpha}}$, so $f=f_{\alpha}$ in $\mathcal{M}$. Since $\dot{f}$ is a $\mathbb{P}_{\alpha}$–name of bounded rank, and no forcing past stage $\alpha$ adds sets of that rank, Theorem 9.2 of Karagila [9] implies that $\dot{f}^{G_{\lambda}}=\dot{f}^{G_{\alpha}}$. Hence $f=f_{\alpha}$ in $\mathcal{M}$.
@@
Remark 5.63 (Countable-support names at limits of cofinality $\omega$).
At limits $\lambda$ with $\operatorname{cf}(\lambda)=\omega$, a hereditarily symmetric name may have countable support that is cofinal in $\lambda$, so stage-bounding arguments do not apply. The iteration machinery of [5] handles such limits using limit-filter $\omega_{1}$-completeness together with the appropriate block/factorization analysis (cf. [5, Rem. 3.9, Thm. 5.5]), and we rely on those limit-stage conclusions.
@@
By equivariance of name interpretation ([8, Lemma 14.37]):
@@
By Lemma 5.6, each orbit package $\mathbb{S}_{[f]}$ used at stage $\alpha$ is $\omega_{1}$-closed (and hence adds no new $\omega$-sequences of ordinals), and the finite-support product $\dot{\mathbb{S}}_{\alpha}$ is again $\omega_{1}$-closed for the same reason. By Lemma 5.16, the relevant stage filter is $\omega_{1}$-complete. Therefore the $\mathsf{DC}$-preservation theorem for symmetric extensions applies (API item (c), cf. [5, Thm. 3.20]), yielding $M_{\alpha+1}\models\mathsf{DC}$.
At limit stages, $\mathsf{DC}$ is preserved by the countable-support symmetric iteration machinery (API item (c), and the limit-filter generation/completeness from API item (d); cf. [5, Thm. 3.20] and [5, Thm. 5.5] for the limit construction).
@@
By Proposition 5.62, $\mathcal{M}\models\mathsf{PP}^{\mathrm{split}}\!\restriction T$, hence $\mathcal{M}\models\mathsf{PP}\!\restriction T$. By Proposition 5.71, $\mathcal{M}\models\mathsf{AC}_{\mathsf{WO}}$. By Lemma 5.70, $\mathcal{M}\models\mathsf{SVC}^{+}(T)$ for $T=\mathcal{P}(S)$, so the hypotheses of the Ryan–Smith localization theorem (Theorem 4.5, i.e., [17, Prop. 3.17] applied with Ryan–Smith’s parameter $S$ taken to be our $T$) hold in $\mathcal{M}$. Therefore $\mathcal{M}\models\mathsf{PP}$.